Rus and dufus each took some money froma piggy bank to buy an ice cream cone but rufus was 24 cent short and dufus was 2 cent short of the price of cone. they decided to pool their resources, but found they still could not afford to buy the cone. how many cents did the ice cream cost?
Answer by ankor@dixie-net.com(22740)

Rus and dufus each took some money from a piggy bank to buy an ice cream cone but rufus was 24 cent short and dufus was 2 cent short of the price of cone. they decided to pool their resources, but found they still could not afford to buy the cone. how many cents did the ice cream cost?
:
Let the cost of the cone = x
then
x - 24 = R's amt
and
x - 2 = D's amt
:
The total amt of both:
(x-24) + (x-2) = (2x - 26)
:
Total amt was not enough to buy a cone, write an inequality
2x - 26 < x
2x - x < 26
x < 26 cents
Let x = 25 cents be the price of a cone
;
Is this true?
25 - 24 = 1 cents for R
and
25 - 2 = 23 cents for D
:
Together they had 24 cents, not enough for a 25 cent cone